Appendix. Picture Gallery Polen, 1892-1945 Patrick Billingsley USA, Contemporary Polen/USA, 1899-1982 France, 1871-1956 Sweden, 1893-1985 USA, 1910- , Contemporary Russia/USA, 1924- , Contemporary Jugoslavia/USA, 1906-1970 Germany, 1777-1855 Russia, 1894-1959 Russia, 1903-1987 France, 1886-1971 Germany, 1832-1903 Yurii Vasilevich Prokhorov USA, 1907-1977 Anatolii Vladimirovich Skorokhod Russia, 1821-1894 USA, 1894-1964